find the range of the function given f(x)=-x^2+3x-5 ; domain={-1,3}
Answer by rmromero(383)


find the range of the function given f(x)=-x^2+3x-5 ; domain={-1,3}
to find the range, we are going to substitute values of x to f(x)=-x^2+3x-5
When x = -1
f(x)= -x^2 + 3x - 5
= -(-1)^2 + 3(-1) - 5
= -1 - 3 - 5
= -9

When x = 3
f(x)= -x^2 + 3x - 5
= -(3)^2 + 3(3) - 5
= -9 + 9 - 5
= -5
Range = {-9, -5}